OpenStudy (anonymous):

a plane traveled 285 miles to Perth and back. The trip there was with the wind. It took 3 hours. The trip back was into the wind. The trip back took 5 hours. What is the speed of the plane in still air? What is the speed of the wind?

OpenStudy (anonymous):

x = speed of the plane y = speed of the wind one way trip is 285/2 = 142.5 miles d = rt d = (x + y)3 (the wind helps you out so you add x and y) d = (x - y)5 (the wind hurts you so you subtract x and y) 142.5 = 3x + 3y 142.5 = 5x - 5y Solve Multiply top by -5 and bottom by 3 so the y's cancel -5(142.5 = 3x + 3y) 3(142.5 = 5x - 5y) do you know how to solve from here?

OpenStudy (anonymous):

yeah. thanks :)